Circular No. 7542

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

XTE J1118+480 J. McClintock, M. Garcia, P. Zhao, N. Caldwell, and E. Falco, Harvard-

Smithsonian Center for Astrophysics, write: “Spectroscopic observations of this x-ray nova were made with the new 6.5-m MMT on Dec. 1 (two spectra) and 4 (six spectra). As expected, the spectra contain both broad Balmer emission lines and the numerous absorption features characteristic of a late-type dwarf. Robust velocity measurements were obtained by crosscorrelating the individual spectra (range 490–650 nm) against the spectra of several bright dwarfs (K5–M2). The eight velocities provide good phase coverage and yield a velocity amplitude K = 698 ± 11 km/s, an orbital period P = 0.1701 ± 0.0003 day (IAUC 7412, 7418), and a time of maximum velocity on 2000 Dec. 1.4776 ± 0.0005 UT. The large value of the mass function, f(M) = 6.0 ± 0.3 M_{⊙}, is rivaled only by that of V404 Cyg. Our Oct. 26 observations (V = 18.8; Whipple Observatory 1.2-m telescope) showed that the nova was near its pre-outburst brightness (IAUC 7390). See also the simultaneous Nov. 30–Dec. 1 photometry given below.”

R. M. Wagner, LBT Observatory; C. B. Foltz, MMT Observatory; S. G. Starrfield, Arizona State University; and P. Hewett, Institute of Astronomy, Cambridge, report: “Thirteen optical spectra (range 410–750 nm; resolution 0.35 nm) of XTE J1118+480 were obtained with the 6.5-m MMT on Nov. 20.40–20.48 and 30.36–30.51 UT. Convolution of the average fluxcalibrated spectrum obtained on Nov. 20 with a V -band filter gives V = 19.0, comparable to its pre-outburst quiescent brightness (IAUC 7390). The spectrum exhibits broad (FWHM ∼ 2400 km/s), double-peaked, Balmer emission lines superposed on the spectrum of an M0–M1V star, similar to other x-ray novae in their quiescent states. Cross-correlation of the individual spectra with three bright, late-type dwarf star spectra yields a sinusoidal variation in radial velocity of the absorption lines with the following orbital parameters: K = 695 ± 16 km/s, P = 0.1699 ± 0.0001 day, T_{0} = 2000 Nov. 20.392 UT ± 0.045 day (corresponding to absolute phase 0.0 or inferior conjunction of the dwarf star). The inferred mass function is 5.9 ± 0.4 M_{⊙}, suggesting that the compact object is a black hole.”

P. Garnavich, University of Notre Dame, monitored XTE J1118+480 via 2-min I-band exposures for a total of 6 hr with the 1.8-m Vatican Advanced Technology Telescope on Nov. 30 and Dec. 1 UT: “The star was near quiescence at V = 18.8 (P. Groot, priv. comm.). The light curve, phased to a period of 4.1 hr (Uemura et al. 2000, PASJ 52, 15), shows a double sinusoidal variation with a peak-to-peak amplitude of 0.15 mag, plus some additional flux at phase 0.55 away from minimum light.”

2000 December 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green